# REST API Testing in DevOps: A Study on an Evolving Healthcare IoT Application

### v1

@@ 4.4 Data Analysis and Interpretation
Table 3 presents analysis metrics, measures, and statistical tests for each RQ, along with its corresponding evaluation aspect. To analyze RQ1 results related to failures, we computed unique failures representing server-side (5XX) and client-side (4XX) errors, accounting for the possibility of multiple occurrences of the same error type. For this purpose, we first grouped the 5XX and 4XX errors separately. We then calculated the similarity between failure responses using the Gestalt approach for string matching proposed by Ratcliff and Metzener [55], implemented in a Python library named difflib [56]. Given that the responses obtained in our study contain data in JSON and HTML formats, we opted for the Gestalt string matching algorithm, considering its application for code comparison [57]. The Gestalt string matching algorithm calculates the similarity between two strings on a range from 0 to 1, where 0 indicates no similarity and 1 denotes a 100% similarity [57, 56]. A threshold of 0.5 (or 50%) serves as a balanced midpoint on this scale. To identify unique failure responses in our study, we set a similarity threshold of 0.5. Responses with a similarity score below 0.5 are considered to exhibit low similarity, while those with a score above this threshold are deemed to have high similarity.
To analyze RQ2 results concerning faults, we followed a semi-automatic approach inspired by Martin-Lopez et al. [27] to discover potential faults. Initially, we grouped unique failure and success responses based on their similarity score, calculated using the Gestalt algorithm with the same threshold set for RQ1. We chose to include the success responses in this process, as our previous work has shown that even successful responses can potentially lead to faults [58]. Next, we created clusters for different groups by parsing the JSON and HTML responses to obtain information about the failures, such as error messages. Lastly, we carefully inspected the resulting list of potential faults and assigned each a unique identifier (ID) to facilitate precise interpretation and reporting of these faults.
@@
To analyze coverage results for RQ3, we used Restats [59] to calculate the coverage of test cases generated by each REST API testing tool. The coverage reported by Restats encompasses different aspects, including paths, operations, status classes, statuses, response types, request types, parameters, and parameter values. For our analysis, we evaluated the overall coverage achieved by each testing tool for all releases.
@@
To analyze RQ4 results related to regressions, we first assembled responses into pairs of two consecutive releases, such as R1-R2, R2-R3, etc. Subsequently, we compared the response status codes of each pair (for example, R1 and R2) to analyze any discrepancy between the status codes of two consecutive releases. If two status codes differed, we further compared the detailed responses of each pair by calculating their similarity using the Gestalt algorithm with the same threshold set for RQ1. Finally, we thoroughly inspected the set of potential regressions and marked each with a unique identifier (ID) to differentiate regressions while reporting results.
@@
In the cost analysis for RQ5, we initially computed the failure and fault detection rates and subsequently calculated the percentages of cost overhead. We utilized Equation 1 to calculate the failure detection rates, and Equation 2 for computing fault detection rates. In these equations, $N_{failures}$ represents the total number of unique failures, $T$ denotes the total number of tests generated by a tool, and $M_{faults}$ symbolizes the total number of unique faults identified from failure results. To compute the cost overhead for the rates corresponding to failures or faults, we use Equation 3. In this equation, we subtract the rates associated with failures or faults from 100. The primary objective is to determine the percentages of failed tests to identify unique failures or faults, thus representing the overhead.
$Rate_{failure}=\left(\frac{N_{failures}}{T}\right)\times 100$ (1)
$Rate_{fault}=\left(\frac{M_{faults}}{N_{failures}}\right)\times 100$ (2)
$Cost=100-Rate_{failure/fault}$ (3)
@@
To statistically compare REST API testing tools for RQ1 and RQ3, we used non-parametric tests adhering to the recommended guidelines by Arcuri and Briand [60]. Specifically, we employed the Wilcoxon test to determine statistically significant differences, with a significance level set to 0.05 (i.e., $\alpha=0.05$). To further analyze which tool performed better, we calculated the effect size using the Vargha-Delaney Â12 measure. For RQ2 and RQ4, we present the results descriptively, focusing on discussions related to potential faults and regressions, similar to existing studies [27, 12].
@@
Table 5: RQ1 results: Unique failures categorized as 4XX and 5XX detected by each tool across all APIs from releases R1 to R14. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
@@
Release 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX
[W]R1 1198 4 1322 7 1209 6 1035 60 913 159 1229 31 169 23 92 27 142 31
[D]R2 394 7 531 4 381 5 569 61 4876 755 217 19 541 79 535 82 608 124
[D]R3 394 11 532 5 381 9 569 59 4876 751 217 13 541 78 535 73 610 119
[W]R4 394 9 532 6 381 5 569 61 4872 746 217 18 541 77 535 74 610 112
[D]R5 394 12 532 4 381 10 569 63 4875 752 217 996 541 79 535 82 610 126
[D]R6 394 10 532 7 380 10 569 61 4876 1633 217 11 541 80 535 79 610 118
[W]R7 394 10 532 6 381 6 569 59 4877 754 217 11 541 78 535 81 610 117
[D]R8 394 6 532 5 381 10 569 64 4878 758 217 15 541 79 535 80 610 117
[D]R9 394 9 532 8 380 10 568 64 4876 2810 217 12 540 83 535 80 610 122
[W]R10 394 9 532 5 381 6 568 61 4877 1156 217 12 541 81 534 82 610 122
[D]R11 394 12 532 7 381 8 568 65 4876 2090 217 22 541 77 534 78 610 128
[D]R12 395 6 533 4 382 3 662 61 5230 4059 218 12 541 86 535 83 797 115
[W]R13 395 3 533 2 382 7 663 58 5281 757 218 18 541 78 535 84 797 118
[M]R14 395 6 533 4 382 4 663 62 5282 1948 218 12 540 80 535 83 797 116
@@
Table 6: Statistical comparison of various testing tools based on the number of failures detected. It includes the p-values derived from the Wilcoxon test and the $\hat{A}_{12}$ values obtained through Vargha-Delaney’s effect size measure.
@@
Figure 4: RQ1 results: Boxplot representing the median number of unique failures detected by each tool across all releases.
Figure 5: RQ2 results: Faults and their frequency of occurrence corresponding to each testing tool. Each fault is color-coded to highlight its uniqueness. The size of the bubble indicates the frequency of occurrence of each fault. Detailed information for these faults is presented in Table 8.
Figure 6: Regressions and their frequency of occurrence corresponding to each testing tool. Each type of regression is color-coded to highlight its uniqueness. The size of the bubble indicates the frequency of occurrence of each regression. Detailed information for all types of regressions is presented in Table 12.

### v2

@@ 4.4 Data Analysis and Interpretation
Table 3 presents analysis metrics, measures, and statistical tests for each RQ, along with its corresponding evaluation aspect. To analyze RQ1 results related to failures, we computed total API failures representing server-side (5XX) errors. We also calculate failure ratios using Equation 1, where $N_{failures}$ represents the total number of failures, $T$ denotes the total number of tests generated by a tool.
To analyze RQ2 results concerning faults, we followed a semi-automatic approach inspired by Martin-Lopez et al. martin2022online to discover potential faults. Initially, we grouped all failure and success responses (string data) based on their similarity score, calculated using the Gestalt algorithm with the same threshold (i.e., 0.5) set for RQ1. We chose to include the success responses in this process, as our previous work has shown that even successful responses can potentially lead to faults sartaj2021testing . To calculate the similarity between responses (string data), we used the Gestalt approach for string matching proposed by Ratcliff and Metzener ratcliff1988pattern , implemented in a Python library named difflib python3difflib . Given that the responses obtained in our study contain data in JSON and HTML formats, we opted for the Gestalt string matching algorithm, considering its application for code comparison tsikerdekis2018persistent ; gabel2010scalable . The Gestalt string matching algorithm calculates the similarity between two strings on a range from 0 to 1, where 0 indicates no similarity and 1 denotes a 100% similarity tsikerdekis2018persistent ; python3difflib . To identify unique responses in our study, we set a threshold of 0.5 (or 50%) as a balanced midpoint on the similarity scale. Responses with a similarity score below this threshold are considered to exhibit low similarity, while those above are deemed highly similar. This threshold is particularly relevant in contexts where structural similarity is important, such as pattern matching and code or data comparisons ratcliff1988pattern ; tsikerdekis2018persistent ; gabel2010scalable . Our choice was inspired by threshold values previously determined through experiments in code similarity research gabel2010scalable . The potential implications of using this threshold include tolerance for minor variations, handling structural similarities, and flexibility in determining what constitutes a unique response. We illustrate the usage of Gestalt string matching in our context with an example shown in Figure 4, where two error responses in JSON format differ only in a single property value. Although the data property in Response 2 contains a slightly longer string, the similarity score computed using the Gestalt algorithm is 0.89, which exceeds the defined threshold and indicates a high similarity. This is primarily due to the Gestalt algorithm’s ability to intuitively compare text, manage minor variations, and handle structural similarities of JSON responses. In another case, when a property such as data in Response 2 contains a significantly longer and differing string compared to Response 1, Gestalt may return a similarity score close to or below the threshold. Therefore, as a follow-up step, we conducted manual inspections on all Gestalt similarity results to ensure the accuracy of the responses’ uniqueness.
After calculating similarity scores, we created clusters for different groups by parsing the JSON and HTML responses to obtain information about failures, such as error messages. Lastly, we carefully inspected the resulting list of potential faults and assigned each a unique identifier (ID) to facilitate the precise interpretation and reporting of these faults. In addition, we calculate fault ratios using Equation 1, where $N_{faults}$ represents the total number of unique faults and $T$ denotes the total number of tests generated by a tool. To further analyze fault detection effectiveness of testing tools over the test suite lifecycle, we calculate the Average Percentage of Faults Detected (APFD), originally introduced by Elbaum et al. elbaum2002test and later applied in a similar context by Martin-Lopez et al. martin2022online .
@@
To analyze coverage results for RQ3, we used Restats corradini2021restats to calculate the coverage (numerical data) of test cases generated by each REST API testing tool. The coverage reported by Restats encompasses different aspects, including paths, operations, status classes, statuses, response types, request types, parameters, and parameter values. For our analysis, we evaluated the overall coverage achieved by each testing tool for all releases. Since we lack access to the source code, we rely on the black-box coverage criteria proposed by Martin-Lopez et al. martin2019test and implemented in Restats.
@@
To analyze RQ4 results related to regressions (string data), we first assembled responses into pairs of two consecutive releases, such as R1-R2, R2-R3, etc. Subsequently, we compared the response status codes of each pair (for example, R1 and R2) to analyze any discrepancy between the status codes of two consecutive releases. If two status codes differed, we further compared the detailed responses of each pair by calculating their similarity using the Gestalt algorithm with the same threshold set for RQ1. Finally, we thoroughly inspected the set of potential regressions and marked each with a unique identifier (ID) to differentiate regressions while reporting results. It is important to note that comparing differences between releases was impossible for us due to the lack of access to API implementations in various releases. Consequently, our regression analysis is based on API responses obtained during test executions across these releases.
@@
In the cost analysis for RQ5, we initially computed the failure and fault detection rates (numerical data) and subsequently calculated the percentages of cost overhead. We utilized Equation 1 to calculate the failure detection rates, and Equation 2 for computing fault detection rates. In these equations, $N_{failures}$ represents the total number of failures, $T$ denotes the total number of tests generated by a tool, and $M_{faults}$ symbolizes the total number of unique faults identified from failure results. To compute the cost overhead for the rates corresponding to failures or faults, we use Equation 3. In this equation, we subtract the rates associated with failures or faults from 100. The primary objective is to determine the percentage of tests that failed to identify failures or faults, thus representing the overhead. In DevOps environments with rapid and frequent releases, it becomes important for practitioners to understand how effectively their testing budget is used, specifically in terms of whether tests contribute to detecting failures or faults. For example, consider a scenario where the entire testing budget (100%) is allocated to execute 300 tests for a specific release. If 12 failures are detected from 300 test cases, a failure detection ratio of 4% is obtained with Equation 1. Subsequently, applying Equation 3, the cost overhead is calculated as 100–4=96%, indicating that 96% of the tests did not reveal any failures. This measurement allows practitioners to estimate testing cost overheads and make better decisions about carefully selecting or prioritizing tests, which is especially useful during continuous delivery cycles in DevOps.
$Ratio_{failure}=\left(\frac{N_{failures}}{T}\right)\times 100$ (1)
$Ratio_{fault}=\left(\frac{M_{faults}}{T}\right)\times 100$ (2)
$Cost=100-Ratio_{failure/fault}$ (3)
@@
To statistically compare REST API testing tools for RQ1 and RQ3, we first assessed (numerical) data normality using the Shapiro-Wilk test shapiro1965analysis , which indicated non-normal distributions with p-values close to 0. As a result, we used non-parametric tests, specifically the Wilcoxon signed-rank test to identify statistically significant differences, and the Vargha-Delaney Â12 effect size measure to analyze which tool performed better, following the recommended guidelines by Arcuri and Briand arcuri2011practical . In addition, we applied Holm’s method holm1979simple to adjust p-values for multiple comparisons, using a significance level of $\alpha=0.05$. The Holm method is a stepwise correction technique that is known for a higher statistical power compared to methods such as Bonferroni correction wagner2021code , which makes it well-suited for our experimental analysis. For RQ2 and RQ4, we present the results descriptively, focusing on discussions related to potential faults and regressions, similar to existing studies martin2022online ; godefroid2020differential .
@@
4.5 Results Discussion
Table 4 presents the number of tests each tool generated for every API. It can be observed that EvoMaster-BB consistently generated tests for every API and produced the greatest number of tests compared to all other tools. In addition, RestTestGen-NET, RESTler-F, RESTler-FL, and Schemathesis-PT were ranked second, and four variants of RESTest were ranked third in terms of the total number of tests generated. From the results, it is noticeable that RestTestGen-MAST did not generate any tests for all APIs, even after successful runs. Given that RestTestGen-MAST is primarily designed to test for security vulnerabilities, it only generates tests upon identifying potential vulnerabilities (as mentioned in Github issue [^2] https://github.com/SeUniVr/RestTestGen/issues/17). This could potentially be attributed to the presence of read-only parameters in the API schema. However, we found that the APIs accessible to us do not contain such parameters. Alternatively, this could be related to the API implementation, which remains a black box to us due to our lack of access to the source code.
In subsequent subsections, we discuss results associated with failures, faults, coverage, regressions, and cost overhead corresponding to each RQ. For a discussion on the results, we refer to Tables 5, 7, 6 and 8 and Figure 5 for RQ1 concerning failures, Tables 9, 10 and 11 and Figures 6 and 7 for RQ2 pertaining to faults, Tables 12, 13 and 14 for RQ3 associated with coverage, Tables 15, 16 and 17 and Figure 8 for RQ4 related to regressions, and Tables 18 and 19 for RQ5 results concerning cost overhead. Note that while discussing the experimental results for each RQ, we maintain a balance between conveying essential information and omitting specific details to prevent disclosing sensitive information that could potentially affect the confidentiality and privacy of Oslo City’s industry partner.
4.5.1 RQ1 Results – REST API Failures
From the results in Table 5, it can be observed that EvoMaster-BB, which generated the most tests, also led to the highest number of API failures compared to other tools. The second highest number of failures was detected in the case of Schemathesis-PT and RestTestGen-NET. This is also evident from Figure 5 that EvoMaster-BB has the highest median number of failures, while Schemathesis-PT and RestTestGen-NET are in the second position. Moreover, the median number of detected failures remains comparable across all variants of RESTest, with RESTler-F and RESTler-FL slightly exceeding the RESTest variants in failure detection. When examining the failure detection ratios, EvoMaster-BB shows lower failure detection ratios compared to Schemathesis-PT and RestTestGen-NET. Among these, Schemathesis-PT demonstrates the highest failure detection ratios, outperforming all other tools. The analysis of the overall results presented in Table 6 shows that EvoMaster maintains the highest total failures, but lower failure detection ratios. This is mainly because EvoMaster generated the highest number of tests, with only a small fraction resulting in failures. Similarly, while RESTler and RestTestGen show comparable total failures, failure detection ratios for RESTler are lower due to the high number of tests generated by its two variants RESTler-F and RESTler-FL. In the case of RESTest, the failure detection ratios are lower compared to all other tools, except for the release R1.
When comparing releases from R1 to R14, it is evident that all tools detected fewer failures in R1 than in subsequent releases. Notably, only the tests generated by RESTest-RT and RESTest-CBT led to identifying a high number of failures in R1, while RESTler-F demonstrated the lowest failure detection. As shown in the overall tool comparison in Table 6, RESTest detected the highest number of failures in R1. Another observation is that tests generated by some tools resulted in a higher number of client-side (4XX) errors in R1. This is especially noticeable with Schemathesis-PT, RESTler-F, RESTler-FL, and RestTestGen-NET, where the ratio of 4XX errors exceeds 90%. From release R2 onward, failure detection by each tool remained relatively consistent with minor variations. Our observations revealed that a significant change in the number of failures from R1 to R2 was primarily due to a shift in the test environment. Release R1, in the staging environment, was transitioned from the prior release in the production environment. This environmental shift induced instability in R1, resulting in a lower detection of failures and higher counts of 4XX errors in the releases from R2 onward.
The results of the statistical comparison of all REST API testing tools, based on their ability to detect failures, are presented in Table 8. The comparative analysis reveals that RESTest-CBT outperformed both RESTest-ART and RESTest-FT, while its failure detection performance was comparable to RESTest-RT because no statistically significant difference was observed between them. While RESTest-ART managed to outperform RESTest-FT, it falls short when compared to RESTest-RT. RESTest-FT, on the other hand, was unable to surpass any of the tools. Notably, EvoMaster-BB demonstrated superior performance over all REST API testing tools. In the case of Schemathesis-PT, it outperformed all variants of RESTest, including RESTest-CBT, RESTest-ART, RESTest-RT, and RESTest-FT, as well as RESTler-F and RESTler-FL. However, no statistically significant difference was observed between Schemathesis-PT and RestTestGen-NET. RESTler-F performed well compared to five tools, i.e., RESTest-CBT, RESTest-ART, RESTest-RT, RESTest-FT, and RESTler-FL, whereas RESTler-FL only outperformed all variants of RESTest. Lastly, RestTestGen-NET performed better than all tools except EvoMaster-BB.
The tests generated using EvoMaster-BB lead to the highest number of failures compared to the other REST API testing tools. The overall statistical comparison also shows that EvoMaster-BB outperformed all other tools, indicating its higher failure detection effectiveness. Furthermore, tests obtained from each tool mostly led to client-side errors, suggesting that these tools often produced requests with invalid data, format, or headers. This observation highlights the need to improve data generation capabilities within tools and ensure robust input validation mechanisms on the API side.
@@ 4.5.1 RQ1 Results – REST API Failures
Table 5: RQ1 results: Total number of API 5XX failures detected by each tool and their ratios across all APIs from releases R1 to R14, derived from all runs. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
@@
Release 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio
[W]R1 218 2.4% 106 1.2% 237 2.6% 80 1.3% 159 0.1% 31 0.3% 23 0.1% 29 0.2% 32 0.2%
[D]R2 564 6.2% 212 2.4% 562 6.2% 91 1.5% 10143 8.1% 3385 27.5% 1678 9.5% 1557 9.8% 3122 14.7%
[D]R3 568 6.2% 213 2.5% 566 6.2% 89 1.5% 10139 8.1% 3379 27.5% 1678 9.5% 1549 9.8% 3120 14.7%
[W]R4 669 7.3% 317 3.7% 666 7.3% 194 3.3% 10580 8.4% 3386 27.5% 1678 9.5% 1550 9.8% 3514 16.5%
[D]R5 673 7.4% 316 3.6% 671 7.4% 197 3.3% 10585 8.4% 4363 35.5% 1680 9.5% 1558 9.8% 3522 16.6%
[D]R6 671 7.4% 318 3.7% 670 7.4% 195 3.3% 11465 9.1% 3379 27.5% 1680 9.5% 1553 9.8% 3517 16.5%
[W]R7 671 7.4% 316 3.6% 667 7.3% 193 3.2% 10583 8.4% 3379 27.5% 1678 9.5% 1557 9.8% 3519 16.5%
[D]R8 666 7.3% 317 3.7% 670 7.4% 198 3.3% 10589 8.4% 3383 27.5% 1679 9.5% 1551 9.8% 3517 16.5%
[D]R9 670 7.4% 320 3.7% 671 7.4% 198 3.3% 12642 10.0% 3378 27.5% 1681 9.5% 1556 9.8% 3520 16.5%
[W]R10 670 7.4% 317 3.7% 667 7.3% 194 3.3% 10989 8.7% 3381 27.5% 1680 9.5% 1555 9.8% 3521 16.5%
[D]R11 671 7.4% 319 3.7% 669 7.3% 199 3.3% 11921 9.5% 3388 27.6% 1678 9.5% 1554 9.8% 3525 16.6%
[D]R12 562 6.2% 212 2.4% 560 6.1% 91 1.5% 13441 10.7% 3379 27.5% 1685 9.5% 1558 9.8% 3118 14.7%
[W]R13 560 6.1% 210 2.4% 564 6.2% 88 1.5% 10148 8.1% 3383 27.5% 1677 9.5% 1558 9.8% 3119 14.7%
[M]R14 563 6.2% 212 2.4% 561 6.2% 92 1.5% 11338 9.0% 3381 27.5% 1680 9.5% 1557 9.8% 3117 14.6%
Table 6: RQ1 results: Overall comparison of tools in terms of total number of API 5XX failures detected by each tool and their ratios across all APIs from releases R1 to R14, derived from all runs. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
RESTest EvoMaster Schemathesis RESTler RestTestGen
Release 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio 5XX Ratio
[W]R1 641 2.0% 159 0.1% 31 0.3% 52 0.2% 32 0.2%
[D]R2 1629 4.0% 10143 8.1% 3385 27.5% 3235 9.6% 3122 14.7%
[D]R3 1636 5.0% 10139 8.1% 3379 27.5% 3227 9.6% 3120 14.7%
[W]R4 1846 5.6% 10580 8.4% 3386 27.5% 3230 9.6% 3514 16.5%
[D]R5 1857 5.7% 10585 8.4% 4363 35.5% 3238 9.6% 3522 16.6%
[D]R6 1854 5.6% 11465 9.1% 3379 27.5% 3206 9.5% 3517 16.5%
[W]R7 1847 5.6% 10583 8.4% 3379 27.5% 3234 9.6% 3519 16.5%
[D]R8 1851 5.6% 10589 8.4% 3383 27.5% 3230 9.6% 3517 16.5%
[D]R9 1859 5.7% 12642 10.0% 3378 27.5% 3237 9.6% 3520 16.5%
[W]R10 1848 5.6% 10989 8.7% 3381 27.5% 3235 9.6% 3521 16.5%
[D]R11 1858 5.7% 11921 9.5% 3388 27.6% 3232 9.6% 3525 16.6%
[D]R12 1425 4.3% 13441 10.7% 3379 27.5% 3243 9.7% 3118 14.7%
[W]R13 1422 4.3% 10148 8.1% 3383 27.5% 3235 9.6% 3119 14.7%
[M]R14 1428 4.3% 11338 9.0% 3381 27.5% 3237 9.6% 3117 14.6%
Table 7: Total number of client-side 4XX errors and their ratios across all APIs from releases R1 to R14, derived from all runs. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
Release 4XX Ratio 4XX Ratio 4XX Ratio 4XX Ratio 4XX Ratio 4XX Ratio 4XX Ratio 4XX Ratio 4XX Ratio
[W]R1 5773 63.4% 5842 67.3% 5772 63.4% 4270 71.7% 109500 86.9% 12175 99.0% 16127 90.9% 15728 99.3% 20932 98.4%
[D]R2 6062 66.6% 6120 70.5% 6064 66.6% 4462 74.9% 60586 48.1% 8837 71.9% 14346 80.8% 14242 89.9% 12918 60.7%
[D]R3 6058 66.5% 6119 70.5% 6060 66.5% 4464 75.0% 60590 48.1% 8843 71.9% 14346 80.8% 14250 90.0% 12920 60.7%
[W]R4 5957 65.4% 6015 69.3% 5960 65.4% 4359 73.2% 60149 47.8% 8836 71.9% 14346 80.8% 14249 90.0% 12526 58.9%
[D]R5 5953 65.4% 6016 69.3% 5955 65.4% 4356 73.2% 60144 47.8% 7859 63.9% 14344 80.8% 14241 89.9% 12518 58.8%
[D]R6 5955 65.4% 6014 69.3% 5956 65.4% 4358 73.2% 59264 47.1% 8843 71.9% 14344 80.8% 14246 89.9% 12523 58.8%
[W]R7 5955 65.4% 6016 69.3% 5959 65.4% 4360 73.2% 60146 47.8% 8843 71.9% 14346 80.8% 14242 89.9% 12521 58.8%
[D]R8 5960 65.4% 6015 69.3% 5956 65.4% 4355 73.1% 60140 47.8% 8839 71.9% 14345 80.8% 14248 89.9% 12523 58.8%
[D]R9 5956 65.4% 6012 69.3% 5955 65.4% 4355 73.1% 58456 46.4% 8844 71.9% 14343 80.8% 14243 89.9% 12520 58.8%
[W]R10 5956 65.4% 6015 69.3% 5959 65.4% 4359 73.2% 60150 47.8% 8841 71.9% 14344 80.8% 14244 89.9% 12519 58.8%
[D]R11 5955 65.4% 6013 69.3% 5957 65.4% 4354 73.1% 60001 47.6% 8834 71.9% 14346 80.8% 14245 89.9% 12515 58.8%
[D]R12 6064 66.6% 6120 70.5% 6066 66.6% 4462 74.9% 59953 47.6% 8843 71.9% 14339 80.8% 14241 89.9% 12922 60.7%
[W]R13 6066 66.6% 6122 70.5% 6062 66.6% 4465 75.0% 60581 48.1% 8839 71.9% 14347 80.9% 14241 89.9% 12921 60.7%
[M]R14 6063 66.6% 6120 70.5% 6065 66.6% 4461 74.9% 60603 48.1% 8841 71.9% 14344 80.8% 14242 89.9% 12923 60.7%
Figure 5: RQ1 results: Boxplot representing the median number of API 5XX failures detected by each tool across all releases and all runs.
@@
Table 8: Statistical comparison of various testing tools based on the number of failures detected across releases R1–R14 and all runs. It includes the p-values derived from the Wilcoxon test and the $\hat{A}_{12}$ values obtained through Vargha-Delaney’s effect size measure.
@@
4.5.2 RQ2 Results – REST APIs Faults
The results in Table 9 present the number of unique potential faults discovered from the results generated by each testing tool, along with their respective fault detection ratios across all releases. In release R1, tests generated by RESTest-CBT and RESTest-RT led to finding the highest number of potential faults with the highest fault detection ratios compared to other releases and tools. From R2 onward, their performance remained relatively consistent. Notably, from R4 to R11, tests generated by EvoMaster-BB and RestTestGen-NET led to detecting the most potential faults, followed by Schemathesis-PT. However, in terms of fault detection ratios, RESTest-FT exhibited the highest, while EvoMaster-BB demonstrated the lowest. Analyzing the overall comparison results in Table 10, it is evident that test cases generated by RESTest led to the highest number of potential faults in most releases. The performance of EvoMaster and RestTestGen was second highest and comparable to each other, followed by Schemathesis and RESTler. When considering fault detection ratios, EvoMaster demonstrated the lowest ratios, while Schemathesis achieved the highest in most releases.
Fault F2 was detected from the successful results that returned a 200 status code. This fault was observed in the Authentication API, where requests containing invalid login information resulted in 200 Ok responses. This implies that the system’s authentication process may not be functioning properly, permitting unauthorized access and thereby posing a considerable security risk. Therefore, F2 could potentially be classified as a security fault.
Fault F6 was found in Measurements API using the tests generated by RESTest-ART, RESTest-CBT, and RESTest-RT. This fault was identified when incorrect user IDs resulted in successful responses with 200 Ok status. We observed this during manual inspection, where some randomly generated IDs by the tools appeared valid within the system but differed from the user ID assigned to us for testing, indicating a potential weakness in the API’s access control mechanism. Similarly, fault F7 was found in two device-related APIs, i.e., Medido and Pilly, and using the tests generated by the four variants of RESTest. In this case, empty/invalid JSON objects for these devices led to 200 Ok responses that contained complete information about the corresponding devices. Furthermore, a similar case was observed for the fault F9 that was identified in Users API using tests generated by RESTest-CBT and RESTest-RT. It was noticed that JSON objects with incorrect property values led to 200 Ok responses, providing all user-relevant records such as tasks/surveys. These faults (i.e., F6, F7, and F9) could be due to missing input validations in the API implementation.
The faults F8, F11, F17, and F18 are associated with errors that originated from the database. Fault F8 was discovered in Measurements API from the tests generated by RESTest-CBT and RESTest-RT. This fault, indicated by 409 Conflict status code, was identified when the API implementation tried to insert measurement records for users with IDs that do not exist in the database. The fault F11 was found in Users API from the status code 500 Internal Server Error and from tests generated by EvoMaster-BB, Schemathesis-PT, and RestTestGen-NET. This fault was identified from a database error when the API implementation tried to insert records with missing fields. Similarly to F8, fault F17 was also found in Measurements API due to 409 Conflict error from tests generated by RESTest-CBT and RESTest-RT. While the status code for both faults is the same, the detailed response for F17 indicated a violation of foreign key constraints in the database. This violation occurred when trying to retrieve measurement records using invalid foreign keys. We have marked this fault as unique and have left it to the API development team to determine if they are the same. For fault F18, identified in the Mobile App API, a 500 Internal Server Error occurred due to long values in the request object. In addition, the error response was accompanied by a complete database query. This fault was detected using the tests generated by EvoMaster-BB, RESTler-FL, RESTler-F, RESTest-FT, and RestTestGen-NET.
The faults F10, F13, F14, F15, and F16 are associated with programming errors or problems related to business logic, categories which align with the predefined classification marculescu2022faults . Fault F10 was discovered due to the 500 Internal Server Error response originating from APIs such as Device-Karie, Device-Medido, Device-Pilly, and Users. This fault occurred when the API implementation attempted to iterate over an object of the String type. Fault F13 was identified due to a 400 Bad Request error in Measurements API. This fault pertains to a type mismatch issue encountered during arithmetic operations. The fault F14 was found due to 500 Internal Server Error response from Reimbursements API. The primary cause of this fault was the conversion of a large value from Python’s Int type to C’s Long type. The fault F15 was located due to a 400 Bad Request error in the Authentication API. This fault occurred during the execution of an array operation on a Boolean-type variable. Fault F16 was discovered due to 500 Internal Server Error response from the Users API. This fault occurred while attempting to access Dict items from an object of the String type. Among these faults, F10 was the most frequently occurring, and tests from all tools led to its detection. In the case of F13, it was only identified from tests generated by RESTest-CBT. F14 was discovered from tests produced by both RESTler-FL and RESTler-F, while F15 was only located from the tests generated by RESTler-FL. In addition, F16 was detected using tests obtained from EvoMaster-BB, RestTestGen-NET, and Schemathesis-PT.
Regarding fault detection over time, Figure 7 presents APFD results for each tool across all tested APIs, excluding the Development API, for which no faults were detected. Analyzing the APFDs indicates that all tools exhibited comparable performance across the entire test suite. For most APIs, almost all tools achieved close to 100% APFD within the first 20% of test executions. In some cases, some tools continued to detect faults up to 60% of the test executions, such as RestTestGen-NET for the Catalog API (Figure 7(n)) and RESTler-F for the Care API (Figure 7(p)). These findings suggest that allocating a testing budget of up to 60% is generally sufficient for REST API testing tools to detect most faults.
Tests from RESTest-CBT and RESTest-RT led to finding the highest number of unique potential faults in different APIs. Overall, 18 potential faults were discovered across all releases from the tests generated by all REST API testing tools. These 18 faults belong to various categories, including database failures, business logic flaws, missing data validation, and security-related issues, fitting within the predefined classification marculescu2022faults . Furthermore, for most tools, the highest average percentage of faults was detected within the first 50% of tests, whereas for some tools, additional faults were detected throughout the entire test suite.